OpenStudy (anonymous):

30 basketballs (15 Spalding, 10 Wilson, and 5 other brand-name balls) are on a basketball court. Barry Wood closes his eyes and arbitrarily picks up a ball from the court. Determine the probability that the ball selected is a Wilson or a Spalding

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

15C1/30C1 =1/2

OpenStudy (anonymous):

sorry that will be 15C1/30C1 + 10C1/30C1 =1/2 +1/3 =5/6